# Complete toric varieties with semisimple automorphism group

### v1

@@ Introduction
In this paper we are interested in the automorphism group of complete toric varieties. Notably, a toric variety $X$ always has at least the acting torus $T\simeq(\mathbb{C}^{*})^{n}$ as a subgroup of $\operatorname{Aut}(X)$, but $X$ may admit additional symmetries beyond this torus. In general, $\operatorname{Aut}(X)$ can be quite complicated, but a classical result of Matsumura and Oort [MO67] shows that it has the structure of a locally algebraic group. In particular, $\operatorname{Aut}^{\circ}(X)$ is itself a connected algebraic group. This fact means we can study $\operatorname{Aut}^{0}(X)$ using the tools of algebraic group theory.
In [LLA22], the authors show that if $X$ is a product of two complete toric varieties, then the automorphism group splits accordingly as a direct product, up to permutation of factors. This raises the the following natural question: does a factorization of the identity component of the automorphism group of a toric variety force a corresponding factorization of the variety?
Seeking conditions under which this question can be answered affirmatively, we obtain our first result in this paper that is useful for decomposing a complete toric variety into products by examining the $1$-skeleton of its fan:
@@
Using this theorem, we find that our original question is answered positively when $\operatorname{Aut}_{X}^{\circ}$ is a semisimple algebraic group. In fact, Demazure proved that the automorphism group of a toric variety with a semisimple automorphism group is of the form $\prod_{i=1}^{k}\mathrm{PGL}_{n_{i}}$ [Dem70]. On the other hand, it is well known that if a toric variety has automorphism group $\mathrm{PGL}_{n+1}$, then it is isomorphic to $\mathbb{P}^{n}$, see Corollary˜4.8. Thus, using Theorem˜1.1, we provide an elementary proof of the following theorem:
@@
This last theorem seems to be well-known by experts. But, to our knowledge, it has no published proof.
Aknowledgements. I would like to thank Giancarlo Lucchini for his guidance on this work. I extend this thanks to Alvaro Liendo for his comments on the writting of this paper. This work was partially supported by ANID via beca Doctorado Nacional 2025 Folio 21251132.
@@ 1.1. Preliminaries on Toric Varieties
A subset $\sigma$ of $N_{\mathbb{R}}$ is a convex poluhedral cone if there exists a finite set $S\subset N_{\mathbb{R}}$ such that
@@
A convex polyhedral cone is pointed if it does not contain any vector subspaces of $N_{\mathbb{R}}$. In this work, we restrict ourselves to pointed cones and will simply refer to them as cones. Let $\sigma\subseteq N_{\mathbb{R}}$ be a cone. We say that $\tau$ is a face of $\sigma$ if it satisfies the following property: for every pair $x,y\in\sigma$ such that $x+y\in\tau$, it follows that $x,y\in\tau$. We now state the definition of a fan, which will allow us to properly reference its properties.
@@
Additionaly, if $\bigcup_{\sigma\in\Sigma}\sigma=V$, we say that the fan is complete.
Let $\Sigma$ and $\Sigma^{\prime}$ be two fans in $N_{\mathbb{R}}$ and $N^{\prime}_{\mathbb{R}}$ respectively. A morphism of fans is a linear transformation $\psi:N_{\mathbb{R}}\to N^{\prime}_{\mathbb{R}}$ that restrics to a homomorphism $\psi^{\prime}:N\to N^{\prime}$ such that, for every $\sigma\in\Sigma$, there exists $\sigma^{\prime}\in\Sigma^{\prime}$ with $\psi(\sigma)\subseteq\sigma^{\prime}$. There is a categorical equivalence between fans and normal toric varieties with toric morphisms, wich can be explored in detail in [Ful93] or [CLS11].
Let $\Sigma$ be a fan in $N_{\mathbb{R}}$. The 1-skeleton of $\Sigma$ is defined as the set $\Sigma(1)$ of $1$-dimesional cones in $\Sigma$. Each such cone will be associated with a minimal generator $\rho\in N$, so we will abusively identify $\rho$ with the cone.
@@ 1.2. Preliminaries on Root Data and Root Systems
For each $\alpha\in R$, we define the endomorphisms $s_{\alpha}$ and $s_{\alpha^{\ast}}$ by:
(1) $s_{\alpha}(x)=x-\langle x,\alpha^{*}\rangle\alpha,\ \forall\alpha\in X\quad\text{and}\quad s^{*}_{\alpha}(y)=y-\langle\alpha,y\rangle\alpha^{*}\ \forall y\in X^{*}$
@@
    For all $\alpha\in R,\langle\alpha,\alpha^{\prime}\rangle=2$.
@@
    For all $\alpha\in R,s_{\alpha}(R)=R$ and $s_{\alpha}(R)=R^{*}$.
In the context of automorphism groups of toric varieties, the spaces $X$ and $X^{*}$ will correspond to the character lattice and co-choracter lattice of a torus.
@@
We define $V\coloneqq\langle R\rangle\otimes_{\mathbb{Z}}\mathbb{R}$, where $\langle R\rangle$ is the subgruop generated by $R$ in $X$. If $R\neq\varnothing$, then $R$ is a root system in $V$ (see [Bou08, Ch. VI, $\S$ 1, Definition 1]). That is, it satisfies the following axioms:
@@
Let $T$ be an algebraic torus and $M$ the character lattice of $T$. If $R\subseteq M$ is a root system in $M_{\mathbb{R}}$, then there exists a unique subset $R^{*}\subseteq N\coloneqq\operatorname{Hom}_{\mathbb{Z}\textrm{-mod}}(M,\mathbb{Z})$ such that $(M,N,R,R^{*})$ is a root datum.
@@
Consequently, there exists a unique set of coroots $R^{*}$ fullfilling the conditions a, b and c. Thus, there is a bijection $R\to R^{*}$, and for each root $\alpha\in R$ the endomorphisms $s_{\alpha}$ and $s_{\alpha}^{*}$ satisfy the properties of Definition˜1.2 (i) and (ii) by property b. ∎
@@
Consider any $\sigma\in\Sigma^{\prime}$. Since $\Sigma$ is complete, there exists a subset $C\subseteq\Sigma$ such that:
$\sigma\subseteq\bigcup_{\tau\in C}\tau\quad\text{with}\quad\sigma\cap\tau\neq\{0\}\quad\text{for all}\quad\tau\in C.$
@@
Thus, $\sigma$ can be expressed as a union. If there exists $\tau\in C$ such that $\sigma\cap\tau=\sigma$, then $\sigma$ is a face of $\tau$ and consequently $\sigma\in\Sigma$ by Definition˜1.1.
@@
Suppose to the contrary that $|\Sigma(1)|\leq n$. Let $\Sigma(1)=\{v_{1},\ldots,v_{m}\}$ with $m\leq n$.
If $m<n$, the generators cannot span the $n$-dimensional space, contradicting completeness.
If $m=n$ and $\{v_{1},\ldots,v_{n}\}$ is linearly dependent, then $\dim(\operatorname{Span}\{v_{1},\ldots,v_{n}\})<n$, leading to the same contradiction as before.
Now consider the case where $m=n$ and $\Sigma(1)$ is linearly independent. We show that $v=-\sum_{i=1}^{n}v_{i}$ belongs to no cone in $\Sigma$. Suppose there exists $\sigma\in\Sigma$ containing $v$. Then $v$ would be a non-negative linear combination of some generators:
@@
Since $\Sigma(1)$ is linearly independent, $1+\lambda_{i}=0$ for all $i$, implying $\lambda_{i}=-1<0$, contradicting $\lambda_{i}\geq 0$. Thus $|\Sigma(1)|>n$. ∎
@@ 3. Demazure Roots
Demazure roots were introduced by Demazure in [Dem70], who showed in that same article that this concept is fundamental for studying the automorphism group of complete toric varieties (we will be more precise about this in Section˜3.2). It thus becomes a crucial object for the development of this article and will be used repeatedly. Let $T$ be a maximal torus in a toric variety $X$, both of dimension $n$. Remember that their character module $M=\operatorname{Hom}(T,\mathbb{G}_{\mathrm{m}})$ and co-character module $N=\operatorname{Hom}(M,\mathbb{Z})$ are then free $\mathbb{Z}$-modules of rank $n$.
@@
3.1. Roots of the proyective space
From now we denote by $I_{n}$ the set $\{1,\dots,n\}$ and by $I_{n}^{0}$ the set $\{0,1,\dots,n\}$. In a suitable basis, the standard fan of $\mathbb{P}^{n}$ is formed by taking all strictly convex cones generated by its rays:
@@ 3.2. Relation Between Roots and Root Data
In this subsection, we connect the root datum of linear algebraic groups with the Demazure roots defined at the beginning of this section.
Let $\Sigma$ be a complete fan, $X$ the corresponding toric variety, and $\mathcal{R}(\Sigma)$ its set of Demazure roots. As we said in the introduction, $\operatorname{Aut}_{X}$ is a locally algebraic group and $\operatorname{Aut}_{X}^{\circ}$ is a connected algebraic group.
Fixing a maximal torus $T\subseteq\operatorname{Aut}_{X}$, we obtain the character module $M\coloneqq\operatorname{Hom}(T,\mathbb{G}_{m})$ and its dual $N\coloneqq\operatorname{Hom}(M,\mathbb{Z})$.
How do Demazure roots of $\Sigma$ relate to $\operatorname{Aut}X$? A partial answer is provided by [LLA22, Theorem 2], which synthesizes key results from Demazure:
@@
Let $\Sigma$ be a complete fan and $X$ its corresponding toric variety. Denote by $R(\Sigma)$ the set of Demazure roots of $\Sigma$. For $\alpha\in R(\Sigma)$, let $\rho_{\alpha}$ be the corresponding cocharacter, and define the rational morphism:
@@
    The rational morphism $a_{\alpha}$ defines a faithful $\mathbb{G}_{\mathrm{a}}$-action on $X$. In particular, the induced morphism $t_{\alpha}:\mathbb{G}_{a}\to\operatorname{Aut}_{X}$ is a closed immersion.
@@
    $t^{-1}t_{\alpha}(s)t=t_{\alpha}\big(\chi^{\alpha}(t)s\big)\quad\forall s\in\mathbb{G}_{a},\forall t\in T.$
@@
    The connected component $\operatorname{Aut}^{\circ}_{X}$ is generated by $T$ and the images of $t_{e}$ for all $e\in\mathcal{R}(\Sigma)$. Consequently, $\operatorname{Aut}_{X}$ is smooth.
@@
By [Spr98, Lemma 7.3.3] and the correspondence between Lie subalgebras and connected subgroups [Hum75, §13.1, Theorem], there exists a bijection between roots $\alpha$ of the adjoint action of $T$ on $\mathfrak{g}$ and closed immersions $\iota_{\alpha}:\mathbb{G}_{a}\to G$ satisfying:
$t\iota_{\alpha}(s)t^{-1}=\iota_{\alpha}(\chi^{\alpha}(t)s)\quad\forall t\in T,\forall s\in\mathbb{G}_{a}.$
@@ 4.1. The case of the projective linear group
$\mathcal{P}(\alpha)\coloneqq\left\{p\in N\mid\langle p,\alpha\rangle=-1\ \text{and}\ \langle p,\alpha^{\prime}\rangle\geq-1,\forall\alpha^{\prime}\in R_{n}\setminus\{\alpha\}\right\}.$
@@
$\left|\left\{\alpha\in R_{n}\mid\langle p,\alpha\rangle=-1\right\}\right|>1.$
Furthermore, Definition˜3.1 implies that for every root $\alpha\in R_{n}$ and every ray $p\in\Sigma_{\mathbb{P}^{n}}(1)$, $\langle p,\alpha\rangle\geq-1$. The following proposition provides an explicit description of $\mathcal{P}(\alpha)$.
@@
For any $\alpha\in R_{n}$, either $\alpha=e_{i}^{*}$ for some $i\in I_{n}$ or $\alpha=e_{ij}^{*}$ for some pair $i,j\in I_{n}$. In both cases, any element $\rho$ in the right-hand set satisfies
@@
Let $\alpha\in R_{n}$. To construct the 1-skeleton of a fan with root system $R_{n}$, we select a ray $\rho_{\alpha}$ from the set $\mathcal{P}(\alpha)$. Fixing this ray forces other roots to share $\rho_{\alpha}$. The set of roots associated to a fixed ray is defined as follows:
@@
$\mathcal{R}(\rho_{\alpha})=\bigcup_{j\in\mathrm{supp}(\rho_{\alpha})}\left\{(\epsilon-a_{k})e_{k}^{*}-\epsilon e_{j}^{*}\mid k\in I_{n}\right\}.$
@@
$|\mathcal{R}(\rho_{\alpha})|=|\mathrm{supp}(\rho_{\alpha})|\cdot(n+1-|\mathrm{supp}(\rho_{\alpha})|).$
@@
for some $j\in\mathrm{supp}(\rho_{\alpha})$ and $k\in I_{n}$. Since $j\in\mathrm{supp}(\rho_{\alpha})$, $a_{j}=\epsilon$, and a direct computation gives $\langle\rho_{\alpha},\alpha^{\prime}\rangle=-\epsilon^{2}=-1$. Thus:
$\mathcal{R}(\rho_{\alpha})\supseteq\bigcup_{j\in\mathrm{supp}(\rho_{\alpha})}\left\{(\epsilon-a_{k})e_{k}^{*}-\epsilon e_{j}^{*}\mid k\in I_{n}\right\}.$
To simplify the reverse inclusion, first assume $\epsilon=1$, so $a_{j}=1$ for all $j\in\mathrm{supp}(\rho_{\alpha})$. The case $\epsilon=-1$ follows analogously. Let $\alpha^{\prime}\in\mathcal{R}_{n}$ satisfy $\langle\rho_{\alpha},\alpha^{\prime}\rangle=-1$. We prove:
(4) $\alpha^{\prime}=(1-a_{k})e_{k}^{*}-e_{j}^{*}\quad\text{for some}\quad j\in\mathrm{supp}(\rho_{\alpha}),k\in I_{n}.$
Case 1: If $\alpha^{\prime}=\delta e_{l}^{*}$ with $\delta\in\{-1,1\}$, then $1=\langle\rho_{\alpha},\alpha^{\prime}\rangle=\delta a_{l}$. Since $\epsilon=1$, $a_{l}=1$ and $\delta=-1$. Thus $\alpha^{\prime}=-e_{l}^{*}$ for some $l\in\mathrm{supp}(\rho_{\alpha})$, satisfying (4) with $j=k=l$.
Case 2: If $\alpha^{\prime}=e_{st}^{*}$ for $s,t\in I_{n}$, then $1=\langle\rho_{\alpha},\alpha^{\prime}\rangle=a_{s}-a_{t}$. With $\epsilon=1$, $a_{k}\in\{0,1\}$, so $a_{s}=0$ and $a_{t}=1$. Thus $\alpha^{\prime}$ satisfies (4) with $j=t\in\mathrm{supp}(\rho_{\alpha})$ and $s\in I_{n}\setminus\mathrm{supp}(\rho_{\alpha})$. As these cover all possible $\alpha^{\prime}$, we conclude:
$\mathcal{R}(\rho_{\alpha})=\bigcup_{j\in\mathrm{supp}(\rho_{\alpha})}\left\{(\epsilon-a_{k})e_{k}^{*}-\epsilon e_{j}^{*}\mid k\in I_{n}\right\}.$
We now prove the cardinality formula. First, we claim that for distinct $j,j^{\prime}\in\mathrm{supp}(\rho_{\alpha})$:
@@
If $k,k^{\prime}\in\mathrm{supp}(\rho_{\alpha})$, then $\epsilon-a_{k}=0$ and $\epsilon-a_{k^{\prime}}=0$, implying:
@@
This implies $j=j^{\prime}$. Now suppose $k\in\mathrm{supp}(\rho_{\alpha})$ and $k^{\prime}\notin\mathrm{supp}(\rho_{\alpha})$. Then $j^{\prime}\neq k^{\prime}$ (since $k^{\prime}\notin\mathrm{supp}(\rho_{\alpha})$). By linear independence of $\{e_{k^{\prime}}^{*},e_{j^{\prime}}^{*}\}$, we must have $j=j^{\prime}$ or $j=k^{\prime}$. If $j=j^{\prime}$, this contradicts our assumption $j\neq j^{\prime}$. If $j=k^{\prime}$, then:
@@
contradicting $a_{k^{\prime}}\in\{0,\epsilon\}$. A symmetric argument applies when $k^{\prime}\in\mathrm{supp}(\rho_{\alpha})$ and $k\notin\mathrm{supp}(\rho_{\alpha})$.
Finally, assume $k,k^{\prime}\notin\mathrm{supp}(\rho_{\alpha})$. By linear independence of the dual basis, possible equalities require:
@@
Moreover, for any fixed $j\in\mathrm{supp}(\rho_{\alpha})$, the set:

### v2

@@ Introduction
In this paper we are interested in the automorphism group of complete toric varieties. Notably, a toric variety $X$ always has at least the acting torus $T\simeq(\mathbb{C}^{*})^{n}$ as a subgroup of $\operatorname{Aut}_{X}$, but $X$ may admit additional symmetries beyond this torus. In general, $\operatorname{Aut}(X)$ can be quite complicated, but a classical result of Matsumura and Oort [MO67] shows that it has the structure of a locally algebraic group. In particular, $\operatorname{Aut}^{\circ}(X)$ is itself a connected algebraic group. This fact means we can study $\operatorname{Aut}^{\circ}(X)$ using the tools of algebraic group theory.
In [LLA22], the authors show that if $X$ is a product of two complete toric varieties, then the automorphism group splits accordingly as a direct product, up to permutation of factors. This raises the following natural question: does a factorization of the identity component of the automorphism group of a toric variety force a corresponding factorization of the variety?
To answer this question affirmatively under specific conditions, we first prove a result useful for decomposing a complete toric variety into products by examining the 1-skeleton of its fan:
@@
Using this theorem, we show that our original question is answered positively when $\operatorname{Aut}_{X}^{\circ}$ is a semisimple algebraic group. In fact, Demazure proved that the automorphism group of a toric variety with a semisimple automorphism group is of the form $\prod_{i=1}^{k}\mathrm{PGL}_{n_{i}}$ [Dem70]. Furthermore, it is well known that if a toric variety has automorphism group $\mathrm{PGL}_{n+1}$, then it is isomorphic to $\mathbb{P}^{n}$, see Corollary˜4.8. Thus, using Theorem˜1.1, we provide an elementary proof of the following theorem:
@@
This last theorem seems to be well-known to experts. A proof of this theorem using Lie theory has been communicated to us by Shintaro Kuroki on [Kur10]. We offer an alternative, elementary algebraic proof.
Acknowledgements. I would like to thank Giancarlo Lucchini for his guidance on this work. I woulf also like to thank to Alvaro Liendo for his comments on the writing of this paper. This work was partially supported by ANID via beca Doctorado Nacional 2025 Folio 21251132.
@@ 1.1. Preliminaries on Toric Varieties
A subset $\sigma$ of $N_{\mathbb{R}}$ is a convex polyhedral cone if there exists a finite set $S\subset N_{\mathbb{R}}$ such that
@@
A convex polyhedral cone is pointed or strictly convex if it does not contain any non-zero vector subspaces of $N_{\mathbb{R}}$. In this work, we restrict ourselves to pointed cones and will simply refer to them as cones. Let $\sigma\subseteq N_{\mathbb{R}}$ be a cone. We say that $\tau$ is a face of $\sigma$ if it satisfies the following property: for every pair $x,y\in\sigma$ such that $x+y\in\tau$, it follows that $x,y\in\tau$. We now state the definition of a fan, which will allow us to properly reference its properties.
@@
Additionally, if $\bigcup_{\sigma\in\Sigma}\sigma=V$, we say that the fan is complete.
Let $\Sigma$ and $\Sigma^{\prime}$ be two fans in $N_{\mathbb{R}}$ and $N^{\prime}_{\mathbb{R}}$ respectively. A morphism of fans is a linear transformation $\psi:N_{\mathbb{R}}\to N^{\prime}_{\mathbb{R}}$ that restricts to a homomorphism $\psi^{\prime}:N\to N^{\prime}$ such that, for every $\sigma\in\Sigma$, there exists $\sigma^{\prime}\in\Sigma^{\prime}$ with $\psi(\sigma)\subseteq\sigma^{\prime}$. There is a categorical equivalence between fans and normal toric varieties with toric morphisms, which can be explored in detail in [Ful93] or [CLS11].
Let $\Sigma$ be a fan in $N_{\mathbb{R}}$. The 1-skeleton of $\Sigma$ is defined as the set $\Sigma(1)$ of $1$-dimensional cones in $\Sigma$. Each such cone will be associated with a minimal generator $\rho\in N$, by abuse of notation, we identify $\rho$ with the cone.
@@ 1.2. Preliminaries on Root Data and Root Systems
For each $\alpha\in R$, we define the endomorphisms $s_{\alpha}$ and $s_{\alpha^{*}}$ by:
(1) $s_{\alpha}(x)=x-\langle x,\alpha^{*}\rangle\alpha,\ \forall x\in X\quad\text{and}\quad s^{*}_{\alpha}(y)=y-\langle\alpha,y\rangle\alpha^{*}\ \forall y\in X^{*}$
@@
    For all $\alpha\in R,\langle\alpha,\alpha^{*}\rangle=2$.
@@
    For all $\alpha\in R,s_{\alpha}(R)=R$ and $s_{\alpha^{*}}(R^{*})=R^{*}$.
In the context of automorphism groups of toric varieties, the spaces $X$ and $X^{*}$ will correspond to the character lattice and cocharacter lattice of a torus.
@@
We define $V\coloneqq\langle R\rangle\otimes_{\mathbb{Z}}\mathbb{R}$, where $\langle R\rangle$ is the subgroup generated by $R$ in $X$. If $R\neq\varnothing$, then $R$ is a root system in $V$ (see [Bou08, Ch. VI, $\S$ 1, Definition 1]). That is, it satisfies the following axioms:
@@
Let $T$ be an algebraic torus and $M$ the character lattice of $T$. If $R\subseteq M$ is a root system in $M_{\mathbb{R}}$, then there exists a unique subset $R^{*}\subseteq N\coloneqq\operatorname{Hom}_{\mathbb{Z}}(M,\mathbb{Z})$ such that $(M,N,R,R^{*})$ is a root datum.
@@
Consequently, there exists a unique set of coroots $R^{*}$ fulfilling the conditions a, b and c. Thus, there is a bijection $R\to R^{*}$, and for each root $\alpha\in R$ the endomorphisms $s_{\alpha}$ and $s_{\alpha}^{*}$ satisfy the properties of Definition˜1.2 (i) and (ii) by property b. ∎
@@
Consider any $\sigma\in\Sigma^{\prime}$. Since $\Sigma$ is a complete fan, its support is the whole space. Let $C\subseteq\Sigma$ be the set of cones $\tau\in\Sigma$ such that $\tau\cap\sigma\neq\{0\}$. Then:
$\sigma\subseteq\bigcup_{\tau\in C}\tau.$
@@
Thus, $\sigma$ is expressed as a union. If there exists $\tau\in C$ such that $\sigma\cap\tau=\sigma$, then $\sigma$ is a face of $\tau$ and consequently $\sigma\in\Sigma$ by Definition˜1.1.
@@
Suppose to the contrary that $|\Sigma(1)|\leq n$. Let $\Sigma(1)=\{v_{1},\ldots,v_{m}\}$ with $m\leq n$. If $m<n$, the generators cannot span the $n$-dimensional space, contradicting completeness. If $m=n$ and $\{v_{1},\ldots,v_{n}\}$ is linearly dependent, then $\dim(\operatorname{Span}\{v_{1},\ldots,v_{n}\})<n$, leading to the same contradiction as before.
Now consider the case where $m=n$ and $\{v_{1},\ldots,v_{n}\}$ is linearly independent. We show that $v=-\sum_{i=1}^{n}v_{i}$ belongs to no cone in $\Sigma$. Suppose there exists $\sigma\in\Sigma$ containing $v$. Then $v$ would be a non-negative linear combination of some generators:
@@
Since the generators are linearly independent, $1+\lambda_{i}=0$ for all $i$, implying $\lambda_{i}=-1<0$, contradicting $\lambda_{i}\geq 0$. Thus $|\Sigma(1)|>n$. ∎
@@ 3. Demazure Roots
Demazure roots were introduced by Demazure in [Dem70], who showed in that same article that this concept is fundamental for studying the automorphism group of complete toric varieties (we will be more precise about this in Section˜3.2). It thus becomes a crucial object for the development of this article and will be used repeatedly. Let $T$ be a maximal torus in a toric variety $X$, both of dimension $n$. Recall that their character module $M=\operatorname{Hom}(T,\mathbb{G}_{\mathrm{m}})$ and cocharacter module $N=\operatorname{Hom}(M,\mathbb{Z})$ are then free $\mathbb{Z}$-modules of rank $n$.
@@
3.1. Roots of the projective space
From now on, we denote by $I_{n}$ the set $\{1,\dots,n\}$ and by $I_{n}^{0}$ the set $\{0,1,\dots,n\}$. In a suitable basis, the standard fan of $\mathbb{P}^{n}$ is formed by taking all strictly convex cones generated by subsets of its rays:
@@ 3.2. Relation Between Roots and Root Data
In this subsection, we connect the root datum of linear algebraic groups with the Demazure roots defined at the beginning of this section. Let $\Sigma$ be a complete fan, $X$ the corresponding toric variety, and $\mathcal{R}(\Sigma)$ its set of Demazure roots. As mentioned in the introduction, $\operatorname{Aut}_{X}$ is a locally algebraic group and $\operatorname{Aut}_{X}^{\circ}$ is a connected algebraic group. Fixing a maximal torus $T\subseteq\operatorname{Aut}_{X}$, we obtain the character module $M\coloneqq\operatorname{Hom}(T,\mathbb{G}_{m})$ and its dual $N\coloneqq\operatorname{Hom}(M,\mathbb{Z})$. How do Demazure roots of $\Sigma$ relate to $\operatorname{Aut}X$? A partial answer is provided by [LLA22, Theorem 2], which synthesizes key results from Demazure:
  1. (1)
@@
Let $\Sigma$ be a complete fan and $X$ its corresponding toric variety. Denote by $R(\Sigma)$ the set of Demazure roots of $\Sigma$. For $\alpha\in\mathcal{R}(\Sigma)$, let $\rho_{\alpha}$ be the corresponding ray, and define the rational morphism:
@@
    The rational morphism $a_{\alpha}$ defines a faithful $\mathbb{G}_{\mathrm{a}}$-action on $X$. In particular, the induced morphism $t_{\alpha}:\mathbb{G}_{\mathrm{a}}\to\operatorname{Aut}_{X}$ is a closed immersion.
@@
    $t^{-1}t_{\alpha}(s)t=t_{\alpha}\big(\chi^{\alpha}(t)s\big)\quad\forall s\in\mathbb{G}_{\mathrm{a}},\forall t\in T.$
@@
    The connected component $\operatorname{Aut}^{\circ}_{X}$ is generated by $T$ and the images of $t_{\alpha}$ for all $\alpha\in\mathcal{R}(\Sigma)$. Consequently, $\operatorname{Aut}_{X}$ is smooth.
@@
By [Spr98, Lemma 7.3.3] and the correspondence between Lie subalgebras and connected subgroups [Hum75, §13.1, Theorem], there exists a bijection between roots $\alpha$ of the adjoint action of $T$ on $\mathfrak{g}$ and closed immersions $\iota_{\alpha}:\mathbb{G}_{\mathrm{a}}\to G$ satisfying:
$t\iota_{\alpha}(s)t^{-1}=\iota_{\alpha}(\chi^{\alpha}(t)s)\quad\forall t\in T,\forall s\in\mathbb{G}_{\mathrm{a}}.$
@@ 4.1. The case of the projective linear group
$\mathcal{P}(\alpha)\coloneqq\left\{p\in N\mid\langle p,\alpha\rangle=-1\ \text{and}\ \langle p,\alpha^{\prime}\rangle\geq-1,\forall\alpha^{\prime}\in\mathcal{R}_{n}\setminus\{\alpha\}\right\}.$
@@
$\left|\left\{\alpha\in\mathcal{R}_{n}\mid\langle p,\alpha\rangle=-1\right\}\right|>1.$
Furthermore, Definition˜3.1 implies that for every root $\alpha\in\mathcal{R}_{n}$ and every ray $p\in\Sigma_{\mathbb{P}^{n}}(1)$, $\langle p,\alpha\rangle\geq-1$. The following proposition provides an explicit description of $\mathcal{P}(\alpha)$.
@@
For any $\alpha\in\mathcal{R}_{n}$, either $\alpha=e_{i}^{*}$ for some $i\in I_{n}$ or $\alpha=e_{ij}^{*}$ for some pair $i,j\in I_{n}$. In either cases, any element $\rho$ in the right-hand set satisfies
@@
Let $\alpha\in\mathcal{R}_{n}$. To construct the 1-skeleton of a fan with root system $\mathcal{R}_{n}$, we select a ray $\rho_{\alpha}$ from the set $\mathcal{P}(\alpha)$. Fixing this ray forces other roots to share $\rho_{\alpha}$. The set of roots associated to a fixed ray is defined as follows:
@@
$\mathcal{R}(\rho_{\alpha})=\bigcup_{j\in\operatorname{supp}(\rho_{\alpha})}\left\{(\epsilon-a_{k})e_{k}^{*}-\epsilon e_{j}^{*}\mid k\in I_{n}\right\}.$
@@
$|\mathcal{R}(\rho_{\alpha})|=|\operatorname{supp}(\rho_{\alpha})|\cdot(n+1-|\operatorname{supp}(\rho_{\alpha})|).$
@@
for some $j\in\operatorname{supp}(\rho_{\alpha})$ and $k\in I_{n}$. Since $j\in\operatorname{supp}(\rho_{\alpha})$, $a_{j}=\epsilon$, and a direct computation gives $\langle\rho_{\alpha},\alpha^{\prime}\rangle=-\epsilon^{2}=-1$. Thus:
$\mathcal{R}(\rho_{\alpha})\supseteq\bigcup_{j\in\operatorname{supp}(\rho_{\alpha})}\left\{(\epsilon-a_{k})e_{k}^{*}-\epsilon e_{j}^{*}\mid k\in I_{n}\right\}.$
To simplify the reverse inclusion, first assume $\epsilon=1$, so $a_{j}=1$ for all $j\in\operatorname{supp}(\rho_{\alpha})$. The case $\epsilon=-1$ follows analogously. Let $\alpha^{\prime}\in\mathcal{R}_{n}$ satisfy $\langle\rho_{\alpha},\alpha^{\prime}\rangle=-1$. We prove:
(4) $\alpha^{\prime}=(1-a_{k})e_{k}^{*}-e_{j}^{*}\quad\text{for some}\quad j\in\operatorname{supp}(\rho_{\alpha}),k\in I_{n}.$
Case 1: If $\alpha^{\prime}=\delta e_{l}^{*}$ with $\delta\in\{-1,1\}$, then $1=\langle\rho_{\alpha},\alpha^{\prime}\rangle=\delta a_{l}$. Since $\epsilon=1$, $a_{l}=1$ and $\delta=-1$. Thus $\alpha^{\prime}=-e_{l}^{*}$ for some $l\in\operatorname{supp}(\rho_{\alpha})$, satisfying (4) with $j=k=l$.
Case 2: If $\alpha^{\prime}=e_{st}^{*}$ for $s,t\in I_{n}$, then $1=\langle\rho_{\alpha},\alpha^{\prime}\rangle=a_{s}-a_{t}$. With $\epsilon=1$, $a_{k}\in\{0,1\}$, so $a_{s}=0$ and $a_{t}=1$. Thus $\alpha^{\prime}$ satisfies (4) with $j=t\in\operatorname{supp}(\rho_{\alpha})$ and $s\in I_{n}\setminus\operatorname{supp}(\rho_{\alpha})$. As these cover all possible $\alpha^{\prime}$, we conclude:
$\mathcal{R}(\rho_{\alpha})=\bigcup_{j\in\operatorname{supp}(\rho_{\alpha})}\left\{(\epsilon-a_{k})e_{k}^{*}-\epsilon e_{j}^{*}\mid k\in I_{n}\right\}.$
We now prove the cardinality formula. First, we claim that for distinct $j,j^{\prime}\in\operatorname{supp}(\rho_{\alpha})$:
@@
If $k,k^{\prime}\in\operatorname{supp}(\rho_{\alpha})$, then $\epsilon-a_{k}=0$ and $\epsilon-a_{k^{\prime}}=0$, implying:
@@
This implies $j=j^{\prime}$. Now suppose $k\in\operatorname{supp}(\rho_{\alpha})$ and $k^{\prime}\notin\operatorname{supp}(\rho_{\alpha})$. Then $j^{\prime}\neq k^{\prime}$ (since $k^{\prime}\notin\operatorname{supp}(\rho_{\alpha})$). By linear independence of $\{e_{k^{\prime}}^{*},e_{j^{\prime}}^{*}\}$, we must have $j=j^{\prime}$ or $j=k^{\prime}$. If $j=j^{\prime}$, this contradicts our assumption $j\neq j^{\prime}$. If $j=k^{\prime}$, then:
@@
contradicting $a_{k^{\prime}}\in\{0,\epsilon\}$. A symmetric argument applies when $k^{\prime}\in\operatorname{supp}(\rho_{\alpha})$ and $k\notin\operatorname{supp}(\rho_{\alpha})$.
Finally, assume $k,k^{\prime}\notin\operatorname{supp}(\rho_{\alpha})$. By linear independence of the dual basis, possible equalities require:
@@
Moreover, for any fixed $j\in\operatorname{supp}(\rho_{\alpha})$, the set: